Given utility function U= where PX = 12 Birr, Birr, PY = 4 Birr and the income of the consumer is, M= 240 Birr. A. Find the utility maximizing combinations of X and Y. B. Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result


1
Expert's answer
2022-03-23T14:03:04-0400

"U= X^{0.5}Y^{0.5}"


"MU_x\u200b=0.5\\frac{Y^{0.5}}{X^{0.5\u200b}}"


"MU_y=0.5 \frac {X^{0.5}}{y^{0.5}}"


"{\frac{MU_x}{p_x}}=\frac {MU_y}{p_y}"

"\frac{Y}{X}= \frac{12}{4}"


"P_xX+ p_yY=M"

y= 3x

"X= \frac{1}{3}Y"

Plug into the below equation;

12X+4Y=240

x=10, y=30

b)

"\frac {\partial U}{\partial x \partial y} =\frac {0.25} {(xy)^{0.5}}"

"MRS x.y = \frac {\partial U} {\partial x \partial y} = \frac {0.25}{(10 \times 30)^{0.5}}=0.015"